You are playing a billiard game with your friends. You noticed that 8-ball spins without rolling. What do you think is the inertia of the 8-ball in angular momentum if it weighs 1.5 lbs and its radius is 1 inch?
1
Expert's answer
2020-04-17T10:31:20-0400
"I=\frac{2}{5}mr^2"

"I=\frac{2}{5}(1.5\cdot 0.453592)(0.0254)^2=1.8\cdot 10^{-4}\ kg\cdot m^2"